OpenStudy (anonymous):

Write a linear factorization of the function. f(x) = x^4 + 36x^2

ganeshie8 (ganeshie8):

check the question once

ganeshie8 (ganeshie8):

it must be f(x) = x4 + 36x2

OpenStudy (anonymous):

yea thats it lol

ganeshie8 (ganeshie8):

ty :) now its easy to factor

ganeshie8 (ganeshie8):

first, pull out the gcf

ganeshie8 (ganeshie8):

x4 + 36x2

OpenStudy (anonymous):

x2?

ganeshie8 (ganeshie8):

yes ! take it out common

ganeshie8 (ganeshie8):

x2(x2 + 36)

ganeshie8 (ganeshie8):

\(x^2(x^2+36) = x^2(x^2 + 6^2)\)

ganeshie8 (ganeshie8):

next use this :- \(a^2+b^2 = (a+bi)(a-bi)\)

OpenStudy (anonymous):

thank you..again :)

ganeshie8 (ganeshie8):

\(x^2(x^2+36) = x^2(x^2 + 6^2)\) \(=x^2(x+6i)(x-6i) \)

ganeshie8 (ganeshie8):

np :)